# Unlocking Efficient Large Inference Models: One-Bit Unrolling Tips the Scales

### v1

@@ 1 Introduction
Large neural networks, including Large Language Models (LLMs), have become essential in nearly every inference task due to their exceptional performance. However, this performance comes at a significant cost: scale. For example, Llama 2 and GPT-4 have $65$ billion and $1.76$ trillion parameters, respectively [1]. These models face challenges related to their size, such as requiring vast amounts of memory for storage and retrieval, consuming substantial power, and being inefficient in communicating over networks or saving for inference. This highlights the need for efficient methods like model quantization and pruning. Quantization involves using fewer bits to store each model parameter, while pruning involves setting some parameter values to zero [2, 3, 4]. One effective way to significantly compress a network through quantization is by applying coarse quantization, or one-bit quantization, which converts parameters into binary values [5, 6].
A recent paper demonstrates that an LLM with just $1.58$-bit or ternary weights in its linear layers can perform as well as a high-resolution model [7]. DeepSeek [8] has demonstrated a remarkable ability to achieve state-of-the-art performance in LLMs while significantly reducing computational resources through innovations like 8-bit quantization. In our paper, we aim to take a step beyond quantization alone and show that leveraging the knowledge of the original problem can significantly enhance computational efficiency and reduce the amount of data required during the training process.
In specific neural networks, sparsity naturally comes from the network’s architecture or underlying physic, eliminating the need for techniques like pruning or manually identifying and zeroing out specific neurons [9]. Compared to traditional neural networks, Deep Unrolling Networks (DUNs) typically have far fewer parameters or are sparse by their nature, as they incorporate domain knowledge through the unrolling process [10] (afforded by the integration of the problem-level reasoning or the model, see Fig. 1). These networks are a form of Learning to Optimize (L2O) methods, specifically structured to solve optimization problems efficiently [11].
Although training is resource-intensive and offline, the goal is to achieve faster online application times. For complex problems like nonconvex optimization, where solutions are hard to find, a well-trained L2O method can outperform traditional approaches, often delivering higher-quality results within similar computational budgets [12, 13]. For instance, L2O can recover sparse signals more accurately and efficiently than LASSO [14, 15, 16].
@@ 1.1 Motivations and Contributions
Quantized DUNs tend to be both sparse and quantized simultaneously. To the best of our knowledge, this paper introduces, for the first time, a one-bit DUN, where the linear components of each layer are reduced to one-bit. This architecture inherently results in a network that is both sparse and binary or ternary in nature. However, a key difference between this approach and the quantization method proposed in [7] is that our network is not truly ternary. Unlike traditional ternary schemes, there is no need to identify and zero out specific neurons. Instead, the sparsity arises from the structure of the DUN, making it purely binary.
To quantize the weights, we employ the Quantization-Aware Training (QAT) approach. By simulating the quantization process during training or fine-tuning, the network can adapt to quantization noise [5]. A key drawback of the QAT scheme is the necessity for pre-training to achieve meaningful results, which can be challenging for large-scale networks [17]. However, this issue can be significantly mitigated if the network is sparse, like DUN, leveraging the problem’s knowledge domain. This is because DUN requires fewer parameters and a smaller dataset for training, as demonstrated in our numerical results.
@@
In this paper, we extend this analysis by investigating sparse quadratic programs, using sparse signal recovery as a representative example, and providing the convergence rate for our one-bit algorithm unrolling. A key distinction of our approach is that our proposed algorithm unrolling transforms each iteration of the algorithm into a layer of the unrolled network. This differs from [12], where the same network (with at least 10 layers) is used for each update process in every iteration.
From a theoretical perspective, to ensure the convergence rate of LISTA, the authors of [14] demonstrated that the weights of the DUN must belong to a specific set, referred to as "good" weights, and proved that this set is non-empty. However, ensuring that this set is non-empty or that binary weights can satisfy these conditions presents significant challenges in our work. To address this issue for one-bit LISTA, we modify the proof framework for the convergence rate, adapting it to establish more relaxed conditions that are achievable even with binary matrices. They assumed a fixed sensing matrix in the recent research [14, 15, 16, 13]. However, in our paper, we analyze the robustness of our algorithm even when the sensing matrix varies.
@@ 3 One-bit Unrolling: Training Process and Solvers
Denote the weights or linear transformations within each layer by $\mathbf{W}^{(k)}=[W_{i,j}^{(k)}]\in\mathbb{R}^{n_{1}\times n_{2}}$ for $k\in[K]$. Our approach in QAT consists of two stages, namely stage \@slowromancapi@ and stage \@slowromancapii@.
@@ 4 Towards Sparse Network with Unrolling Algorithm
In this approach, unlike traditional sparsification techniques, sparsity is naturally embedded in the model through predefined zero elements in the weight matrix. This sparsity stems from the iterative update process of the classical algorithm that the network unrolls, where the weight matrix is inherently sparse. Thus, the network structure is not artificially pruned but instead mirrors the sparsity present in the underlying algorithm. By considering this sparsity pattern coming from the nature of the problem, one can hope for reduced complexity in a large-scale regime. It is important to note that this type of sparsification is more computationally efficient compared to the approaches utilized in sparsifying the LLMs since the locations of zero links in the weight matrices of each layer are predefined in our model and there is no need to update the locations during the training.
@@
where $\mathbf{W}_{k}\in\{-\lambda,\lambda\}^{v\times p}$ for $k\in[K]$ are learnable parameters. This block structure ensures that each part of the model interacts only with the relevant input, maintaining the sparsity of the system. The interesting aspect of this model is that as the number of block partitions increases, the sparsity of the network also increases significantly. In some problems, the sensing matrix $\mathbf{A}$ itself is sparse. In such cases, even certain elements within the partitions are independent of one another, generating their outputs based solely on these corresponding elements. These models can further sparsify the weight matrix, making it increasingly sparse.
@@ 5 Generalization Ability
Unlike a set of “good" weights in Definition 1, it is relatively straightforward to demonstrate that a binary weight exists such that $\alpha_{\phi}<1$. To illustrate this, based on our quantization scheme in Section 3, we can express all weights as $\mathbf{W}_{k}=\lambda\mathbf{B}_{k}$, where $\mathbf{B}_{k}\in\{-1,1\}^{n\times n}$ for all $k\in[K]$. Then, we can write
@@
Define $h=\sup_{k\in[K]}\|\mathbf{B}_{k}\|\|\mathbf{Q}\|$. By appropriately selecting values for $\delta$ and $\lambda=\mathcal{O}(h^{-1})$, we can ensure that $\delta+\lambda h<1$. It is important to note that this choice of $\delta$ and $\lambda$ is specifically designed for the worst-case scenario. In Appendix B, we provide further details on how to systematically enforce the constraint $\alpha_{\phi}<1$ during the learning dynamics.
Similar to the ST update process, in the case of HT, our approach involves incorporating $0<\delta\leq 1$ and focuses on analyzing the spectral norm of the form $\left\|\delta\mathbf{I}-\mathbf{W}^{\top}_{\mathcal{S},k}\mathbf{Q}_{\mathcal{S}}\right\|$ for all $k\in[K]$. In Appendices C, D.2, and E.2, we show that the parameter $\alpha_{\phi}=\sup_{k\in[K]}\|\delta\mathbf{I}-\mathbf{W}^{\top}_{\mathcal{S},k}\mathbf{Q}_{\mathcal{S}}\|$ plays a crucial role in all algorithm properties of the HT update process. Regarding order complexity, the impact of $\alpha_{\phi}$ on all algorithm properties in this case is equivalent to the ST update process discussed in Table 1, and as such, we do not repeat it here. Therefore, to guarantee generalization ability for the HT update process, it is required that $\alpha_{\phi}<1$. Further details on how to systematically enforce the constraint $\alpha_{\phi}<1$ during the learning dynamics can be found in Appendix C.
@@
The proof of the aforementioned theorem is straightforward and follows closely to the approach used in the proof of Theorem 5.1. in [12].
@@ Appendix A Convergence Analysis Based on The Set of Good Weights in Definition 1
Before providing the convergence result, we define the residual error as
@@
To generalize (39) for the whole dataset, we take the supremum over $\mathbf{x}^{\mathrm{opt}}(\mathbf{Q},\mathbf{b})$ and by $|\mathcal{S}|=s$ we can write
@@ E.1 Soft-Thresholding
For the scaled binary weights, i.e, $\mathbf{W}_{i}=\lambda\mathbf{B}_{i}$ for $i\in[k]$, we have
@@
According to the order of $c(k-1)$, and by taking supremum w.r.t $\mathbf{Q}$, the results provided in Table 1 can be verified.

### v2

@@ 1 Introduction
Large neural networks, including Large Language Models (LLMs), have become essential in nearly every inference task due to their exceptional performance. However, this performance comes at a significant cost: scale. For example, Llama 2 and GPT-4 have $65$ billion and $1.76$ trillion parameters, respectively [1]. These models face challenges related to their size, such as requiring vast amounts of memory for storage and retrieval, consuming substantial power, and being inefficient in communicating over networks or storage for inference. This highlights the need for efficient methods like model quantization and pruning. Quantization involves using fewer bits to store each model parameter, while pruning involves setting some parameter values to zero [2, 3, 4]. One effective way to significantly compress a network through quantization is by applying coarse quantization, or one-bit quantization, which converts parameters into binary values [5, 6].
A recent paper demonstrates that an LLM with just $1.58$-bit or ternary weights in its linear layers can perform as well as a high-resolution model [7]. DeepSeek [8] has demonstrated a remarkable ability to achieve state-of-the-art performance in LLMs while significantly reducing computational resource needs through innovative use of $8$-bit quantization. In our paper, we take a step beyond quantization alone and show that leveraging the knowledge of the original problem, including models and domain structure, can significantly enhance computational efficiency and reduce the amount of data required during the training process. The associated process, called deep unrolling, leads to a network sparsity that naturally comes from the underlying domain knowledge, eliminating the need for techniques like pruning or manually identifying and zeroing out specific neurons [9]. Compared to traditional neural networks, Deep Unrolling Networks (DUNs) typically have far fewer parameters and are sparse by nature [10] (afforded by the integration of the problem-level reasoning or the model, see Fig. 1). These networks are a form of Learning to Optimize (L2O) models, specifically structured to solve optimization and inference problems efficiently [11].
Although training is often resource-intensive and offline, a worthy goal is to achieve faster online application times. For complex problems like nonconvex optimization, where solutions are hard to find, a well-trained L2O method can outperform traditional approaches, often delivering higher-quality results within similar computational budgets [12, 13]. For instance, L2O can recover sparse signals more accurately and efficiently than LASSO [14, 15, 16].
@@ 1.1 Motivations and Contributions
Quantized DUNs tend to be both sparse and quantized simultaneously. To the best of our knowledge, this paper introduces, for the first time, a one-bit DUN, where the linear components of each layer are reduced to one-bit. This architecture inherently results in a network that is both sparse and binary (or ternary) in nature. However, a key difference between this approach and the quantization method proposed in [7] is that our network is not truly ternary. Unlike traditional ternary schemes, there is no need to identify and zero out specific neurons. Instead, the sparsity arises from the structure of the DUN and is a priori known, making it purely binary.
To quantize the weights, we employ the Quantization-Aware Training (QAT) approach. By simulating the quantization process during training or fine-tuning, the network can adapt to quantization noise [5]. A key drawback of the QAT scheme is the necessity for pre-training to achieve meaningful results, which can be challenging for large-scale networks [17]. However, this issue can be significantly mitigated if the network is sparse, like DUN, leveraging the problem’s domain knowledge. This requires a much smaller dataset for training compute footprint, as demonstrated in our numerical results.
@@
In this paper, we extend this analysis by investigating sparse quadratic programs, using sparse signal recovery as a representative example, and providing the convergence rate for our one-bit algorithm unrolling. A key distinction of our approach is that our proposed algorithm unrolling transforms each iteration of the algorithm into a layer of the unrolled network. This differs from [12], where the same network (with at least $10$ layers) is used for each update process at every iteration.
From a theoretical perspective, to ensure the convergence rate of LISTA, the authors of [14] demonstrated that the weights of the DUN must belong to a specific set, referred to as “good" weights, and proved that this set is non-empty. However, ensuring that this set is non-empty or that binary weights can satisfy these conditions presents significant challenges in our work. To address this issue for one-bit LISTA, we modify the proof framework for the convergence rate, adapting it to establish more relaxed conditions that are achievable even with binary matrices. Recent works assume a fixed sensing matrix [14, 15, 16, 13]. However, in our paper, we analyze the robustness of our algorithm even when the sensing matrix varies.
@@ 3 One-bit Unrolling: Training Process and Solvers
Denote the weights or linear transformations within each layer by $\mathbf{W}^{(k)}=[W_{i,j}^{(k)}]\in\mathbb{R}^{n_{1}\times n_{2}}$ for $k\in[K]$. Our approach in QAT consists of two stages, discussed below.
@@ 4 Towards Sparse Network with Unrolling Algorithm
In this approach, unlike traditional sparsification techniques, sparsity is naturally embedded in the model through predefined zero elements in the weight matrix. This sparsity stems from the iterative update process of the classical algorithm that the network unrolls, where the weight matrix is inherently sparse. Thus, the network structure is not artificially pruned but instead mirrors the sparsity present in the underlying problem. This leads to a reduced complexity that benefits adoption in large-scale regimes. It is important to note that this type of sparsification is more computationally efficient compared to the approaches utilized in sparsifying the LLMs since the locations of zero links in the weight matrices of each layer are predefined in our model and there is no need to update the locations during the training.
@@
where $\mathbf{W}_{k}\in\{-\lambda,\lambda\}^{v\times p}$ for $k\in[K]$ are learnable parameters. This block structure ensures that each part of the model interacts only with the relevant input, maintaining the sparsity of the system. The interesting aspect of this model is that as the number of block partitions increases, the sparsity of the network also increases significantly. In some problems, the sensing matrix $\mathbf{A}$ itself is sparse. In such cases, even certain elements within the partitions are independent of one another, generating their outputs based solely on these corresponding elements. These relations, or lack thereof, help make the inference model increasingly sparse.
@@ 5 Generalization Ability
Unlike the set of “good" weights in Definition 1, it is relatively straightforward to demonstrate that a binary network weight exists such that $\alpha_{\phi}<1$. To illustrate this, based on our quantization scheme in Section 3, we can express all weights as $\mathbf{W}_{k}=\lambda\mathbf{B}_{k}$, where $\mathbf{B}_{k}\in\{-1,1\}^{n\times n}$ for all $k\in[K]$. Then, we can write
@@
Define $h=\sup_{k\in[K]}\|\mathbf{B}_{k}\|\|\mathbf{Q}\|$. By appropriately selecting values for $\delta$ and $\lambda=\mathcal{O}(h^{-1})$, we can ensure that $\delta+\lambda h<1$. It is important to note that this choice of $\delta$ and $\lambda$ is specifically designed for the worst-case scenario. In Appendix B, we provide further details on how to systematically enforce the constraint $\alpha_{\phi}<1$ during the learning process.
Similar to the ST update process, in the case of HT, our approach involves incorporating $0<\delta\leq 1$ and focuses on analyzing the spectral norm of the form $\left\|\delta\mathbf{I}-\mathbf{W}^{\top}_{\mathcal{S},k}\mathbf{Q}_{\mathcal{S}}\right\|$ for all $k\in[K]$. In Appendices C, D.2, and E.2, we show that the parameter $\alpha_{\phi}=\sup_{k\in[K]}\|\delta\mathbf{I}-\mathbf{W}^{\top}_{\mathcal{S},k}\mathbf{Q}_{\mathcal{S}}\|$ plays a crucial role in all algorithm properties of the HT update process. Regarding order complexity, the impact of $\alpha_{\phi}$ on all algorithm properties in this case is equivalent to the ST update process discussed in Table 1, and as such, we do not repeat it here. Therefore, to guarantee generalization ability for the HT update process, it is required that $\alpha_{\phi}<1$. Further details on how to systematically enforce the constraint $\alpha_{\phi}<1$ during the learning process can be found in Appendix C.
@@
The proof of the aforementioned theorem is straightforward and follows closely to the approach used in the proof of Theorem 5.1 in [12].
@@ Appendix A Convergence Analysis Based on the Set of Good Weights in Definition 1
Before providing the convergence result, we define the residual error as follows
@@
To generalize (39) for the whole dataset, we take the supremum over $\mathbf{x}^{\mathrm{opt}}(\mathbf{Q},\mathbf{b})$, and by $|\mathcal{S}|=s$ we can write
@@ E.1 Soft-Thresholding
For the scaled binary weights, i.e., $\mathbf{W}_{i}=\lambda\mathbf{B}_{i}$ for $i\in[k]$, we have
@@
According to the order of $c(k-1)$, and by taking supremum w.r.t. $\mathbf{Q}$, the results provided in Table 1 can be verified.